OpenStudy (anonymous):

What is the sum of a 12-term arithmetic sequence where the last term is 13 and the common difference is -10?

OpenStudy (anonymous):

Here \(n = 12\). \(l = 13\) and \(d = -10). Firstly find out first term by using: \[l = a + (n-1)d\]

OpenStudy (anonymous):

* \(d = -10\)

OpenStudy (anonymous):

\[l = a + (n-1)d\]

OpenStudy (anonymous):

So.. 13 = a +(12-1)-10 13 = a+(11)-10 13= a+1 -1 a=12 ???

OpenStudy (anonymous):

\[S_n=\frac{ n }{ 2 }[2a+(n-1)d]\]

OpenStudy (anonymous):

use the formula waterineyes gave you to find a, since you were already given the values of n and d in the question... \[l=a+(n-1)d\] \[13=a +(12-1) \times -10\] so find a and put into the formula...

OpenStudy (anonymous):

So, 13=12+(12-1)x-10 13= 12+(11)x-10 13=23x-10 +10 23=23x x=1 O.o Now, I'm superr lost lol

OpenStudy (anonymous):

\(\bf d=-10\) --> Common difference \(\bf a_{13}= a + 12d = 13\) Notice, by writing term 13 in terms of 'a' and 'd', we can now solve for 'a' by plugging in -10 for d. \[\bf a+12d=13 \rightarrow a+12(-10)=13 \implies a = 133\]So now we have the first term, 'a', which means we can now use the summation formula to find the sum of the first 13 terms:\[\bf S _{n}=\frac{ n }{ 2 }[2a+(n-1)d]\]Plug in 133 = a, -10 = d, and 13 = n, and evaluate. Can you do that? @Imckeex3

OpenStudy (anonymous):

I'm doing something wrong.. sn = 13/2[2(133)+(13-1)-10] 13/2[266+12-10] 13/2(268)...

OpenStudy (anonymous):

it is supposed to be 12/2 not 13/2......

OpenStudy (anonymous):

\[\bf S_{13} = \frac{ 13 }{ 2 }[2(133)+(13-1)(-10)]=\frac{ 13 }{ 2 }[266+(-120)]=\frac{ 13 }{ 2 }[146]=949\] That's what I got. @lmckeex3

OpenStudy (anonymous):

S(n)=Sum of the n terms in the sequence n=number of terms in the sequence = 12 a=the 1st term in the sequence d=common difference

OpenStudy (anonymous):

Oh sorry I did 13, I forgot that they only wanted the sum of the first 12 lol.

OpenStudy (anonymous):

Just change 13 to 12 and you should get an answer of 936.

OpenStudy (anonymous):

The choices are wrong.

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

there's a mistake....a is supposed to be 123 not 133....

OpenStudy (anonymous):

omg lol

OpenStudy (anonymous):

if you make a the subject of the formula i gave above using the formula waterineyes gave, you get a to be 123... \[13=a + (12-1) \times -10\]

OpenStudy (anonymous):

if you correct the mistake, what do you get now??

OpenStudy (anonymous):

by x.. do you mean *?

OpenStudy (anonymous):

yes i mean multiplication....

OpenStudy (anonymous):

ddnt we say that 13 was actually 12? I'm assuming by x you mean *? so 12=123+(12-1)x-10 12 = 123+11-10 12= 134 * -10 12 = -1340 I'm just going to guess lol. I've completed the quiz. This is the last question.

OpenStudy (anonymous):

\[S _{12} = \frac{ 12 }{ 2 }[2(123) + (12-1) \times -10]\]

OpenStudy (anonymous):

do you see it now??

OpenStudy (anonymous):

6[(246)+(11)(-10) 816 I wasn't using pemdas correctly. THANK YOU!!!!

OpenStudy (anonymous):

you are welcome....(:

OpenStudy (anonymous):

You have found a wrong there.. Sorry, I am late.. \[13 = a + (11)(-10) \implies a = 13 + 110 = a = 123\]

OpenStudy (anonymous):

Here when you find \(a\) and \(l\) here, why you go for that long formula?? Just use: \[Sum = \frac{n}{2}(a + l)\]

OpenStudy (anonymous):

^^ That would have been helpful an hour ago lol.

OpenStudy (anonymous):

\[S_{12} = \frac{12}{2}(123 + 13) = ??\]

OpenStudy (anonymous):

I think this would have been helpful 13 hours ago, when you got offline.. :P

OpenStudy (anonymous):

lol. :x I fell asleep. Thanks to everyone for their help tho. I got 100.